# revtex4-2Repair the float package

### v1

@@ Abstract
Rotation symmetric bosonic codes are are an attractive encoding for qubits into oscillator degrees of freedom, particularly in superconducting qubit experiments. While these codes can tolerate considerable loss and dephasing, they will need to be combined with higher level codes to achieve large-scale devices. We investigate concatenating these codes with the planar code in a measurement-based scheme for fault-tolerant quantum computation. We focus on binomial codes as the base level encoding, and estimate break-even points for such encodings under loss for various types of measurement protocol. These codes are more resistant to photon loss errors, but require both higher mean photon numbers and higher phase resolution for gate operations and measurements. We find that it is necessary to implement adaptive phase measurements, maximum likelihood quantum state inference, and weighted minimum weight decoding to obtain good performance for a planar code using binomial code qubits.
@@ I Introduction
We obtain threshold values for varying orders of discrete rotational symmetry. We find that whilst increasing the order of discrete rotational symmetry generally improves the threshold value, thresholds for binomial encoded qubits tend to be lower than for the trivial Fock space encoding when using the most naive measurement and qubit state inference techniques. By using more sophisticated measurements and state inference methods as well as incorporating the soft information from measurements into decoding we are able to find schemes that beat the trivial encoding when qubit state measurement error for the trivial encoding is not too small.
The competition between the bosonic codes and the trivial encoding arises because the rotationally symmetric codes achieve tolerance to loss errors at the expense of increased Fock number. In comparison, the trivial encoding has no intrinsic robustness to loss but has a very low average number. We find that optimal measurement and state inference is required for the rotational symmetric codes to outperform the trivial encoding. Finally, we examine the sub-threshold performance of binomial codes and find that below $50\%$ of threshold the loss tolerance of a binomial code is more beneficial and codes with increased rotational symmetry perform better than the trivial code and better than codes with reduced symmetry.
This paper is structured as follows. In Sec. II we review rotation symmetric bosonic codes and measurement based quantum computation. In Sec. III we introduce the measurement models and qubit state inference techniques used to realise the computational scheme. Sec. III.4 then quantifies the performance of these methods. In Sec. III.5 we introduce methods used for the 2D planar code, including an X basis Pauli twirling approximation and Dijkstra’s decoding algorithm. Sec. III.6 compares the code thresholds obtained using these techniques, as well as examining the subthreshold scaling of the code. Finally, in Sec. IV we review our findings and highlight the main conclusions.
@@ II.1 Rotation Symmetric Bosonic Codes
Indeed, $\hat{R}_{2N}=\hat{Z}_{N}.$ The choice of the logical Z operator determines the form of the logical codewords as finite rotated superpositions of a “primitive” state $|\theta\rangle$, where different types of rotation symmetric codes, such as cat and binomial codes, differ only in the choice of primitive. Explicitly, for $\mu\in\{0,1\}$
@@ II.1.1 RSB $C_{Z}$ Gate
where modes $i,j$ involve RSB qubits with discrete rotational symmetry of order $N,M$ respectively. The gate time is chosen such that $\Omega t_{\rm gate}=\pi/NM$. In all that follows we only consider CROT gates between identical encodings, so $N=M.$
@@ II.3 Measurement Based Quantum Computation
In this section we will first describe some preliminary notation used for cluster states, before detailing the construction of 1D cluster states. We will then explain how 1D cluster states can be entangled to realise the 2D surface code.
@@ II.3.1 Cluster States
In order to construct the 2D surface code in Fig. 2, we start with parallel 1D cluster states, prepared as before and entangle every odd (primal) qubit of each 1D cluster with the corresponding primal qubit of its neighbouring 1D cluster state, via a dual qubit prepared in the |+ket|$\rangle$| + ⟩ state. Every second qubit of the 1D cluster state is also a dual qubit. The resulting cluster state is depicted in Fig. 4. Measuring every qubit in the X basis realises a measurement of parity checks of given by a product of $X$ on the either qubits around either a primal or dual plaquette. A primal (dual) plaquette stabilizer checks for $Z$ type errors on that plaquette and for $X$ type errors on neighbouring dual (primal) qubits. More explicitly, consider the stabilisers of the cluster state. For a plaquette $p$ of the cluster state, the associated stabiliser is $S_{p}=\prod_{e\in p}X_{e}$, where $e$ denote the primal qubits represented by the edges of the given plaquette as in Fig. 3. To teleport logical information across the 2D foliated state, we measure both primal and dual qubits in the X basis. This also provides stabiliser outcomes, from which we extract the error syndrome.
@@ III.1.4 Measurement Error
Figure 5: Measurement Error vs $\bar{n}$ for binomial code with different $N$ values using heterodyne measurement.
@@
Figure 6: Measurement Error vs $\bar{n}$ for binomial code with different $N$ values using AHD measurement.
@@ III.4.1 Numerical Performance of Phase Measurements
We performed simulations of the minimal instance of the binomial encoded 1D cluster state scheme with a 3-qubit cluster state. Fig. 8 below quantifies the performance of the different measurement schemes. We see that both canonical phase and AHD measurement schemes significantly outperform heterodyne. Furthermore, there does not appear to be a notable difference between the performance of AHD and canonical phase measurements. This indicates that AHD is a good alternative to the optimal canonical phase measurement.
@@ III.5.1 X-Basis Pauli Twirl
III.5.2 Dijkstra’s Algorithm Decoder
@@
During the MWPM algorithm, a complete graph is created from nodes of the syndrome graph. Distances are then calculated between all possible combinations of node pairs. If, for a given node pair, edge weights of the syndrome graph are calculated using soft measurement outcomes, distances cannot be precomputed and must be evaluated using Dijkstra’s algorithm. This means that for each node pair in the complete graph, Dijkstra’s algorithm must be run again. In contrast, if edge weights for a node pair are given fixed unit weight, Dijkstra’s algorithm need only be run a single time. We use a hybrid model in which for nodes separated by Manhattan distances greater than $\log_{2}(L)$, edge weights are calculated using observed hard measurement outcomes, whilst for nodes pairs for which $d_{Manhattan}\leq\log_{2}(L)$, edge weights are calculated using soft measurement outcomes. This provides the advantage of exploiting soft information for a more accurate syndrome decoder, whilst avoiding the long runtime of exhaustively calculating shortest paths between every possible node pair combination.
@@
Finally, we note that we benchmark the performance of the binomial code against the trivial Fock space encoding $|0\rangle_{L}=|0\rangle,|1\rangle_{L}=|1\rangle$, as well as binomial codes of different orders of discrete rotational symmetry.
@@ III.5.4 Overview of Numerical Model
Error correction: We begin by measuring logical Z before error correction. We choose one of the smooth boundaries. If an odd number of boundary qubits have errors, the logical Z measurement is $-1$, otherwise it is $+1$. Error correction proceeds as in Sec. II.3. Stabiliser measurements are modelled by, for each plaquette, multiplying together the stored values of all primal qubits comprising the plaquette. If a plaquette has an odd number of qubits with errors, there will be an odd number of $-1$ values and so the result of the multiplication will be $-1$. Otherwise, it will be $+1$. Stabilisers with $-1$ measurement outcomes form the error syndrome. The BlossomV implementation of the MWPM algorithm is then used to pair up nodes in the error syndrome. As we are modelling the planar code, there can be an odd number of $-1$ stabiliser outcomes, so it is also possible to pair syndrome nodes with the boundary. If the modified Dijkstra’s algorithm decoder is being used, it is used as described in Sec. III.5.2 to calculate weights between pairs of syndrome nodes to feed into MWPM. Otherwise, each edge of the lattice is assigned weight one. The now paired syndrome nodes are used to determine if a logical error has occurred. We pick the same smooth boundary we used for our initial logical Z measurement. For each of the paired syndrome nodes, we check if the path between them crosses our chosen smooth boundary. If an odd number of crosses occur out of all pairs, the logical Z measurement is $-1$, otherwise it is $+1.$ We compare the logical Z measurement before correction to the new logical Z measurement. If they disagree, a logical error has occurred.
@@ III.6.1 Threshold curves
In this section we will refer to codes using heterodyne phase measurement and the binning algorithm QSI as Method 1 Codes, and codes using AHD measurement and local ML QSI as Method 2 Codes.
Fig. 10 shows the threshold for a range of binomial codes against photon loss $\gamma=\kappa t_{\rm gate}$. Binomial codes with varying values of $N$ are compared to the trivial encoding. The data point for a given value of $N$ and $K$ is plotted at the effective measurement error that is determine for this code in section III. This allows a comparison of the threshold for the trivial encoding with a given qubit measurement error. In this initial example, Method 1 codes are used. The most striking feature of this plot is that the trivial encoding significantly outperforms the binomial codes, for all values of $N.$ Furthermore, increasing $N$ does not have an advantageous effect on the threshold, as might be expected. Finally, varying the binomial code parameter $K,$ and therefore varying the effective measurement error (for a code of fixed $N$) also does not appear to significantly change the threshold.
@@
These effects are strongly suggested by Fig. 11, which plots the same data as Fig. 10, except against $\gamma\bar{n}$ rather than $\gamma$. This effectively normalises the data against mean photon number. We see that binomial codes of higher $N$ do better by this measure, and furthermore that binomial codes outperform the trivial encoding. Thus if we look at the thresholds against the probability of a single photon emission occuring $\gamma\bar{n}$ rather than against the probability for each photon to be lost $\gamma$ thne indeed there is increased tolerance to photon loss errors for the higher $N$ codes, although it is the original thresholds measured against $\gamma$ that matter in practice.
@@
Figure 14: Threshold for photon loss errors $\gamma$ with Dijkstra enhanced MWPM decoder. The trivial encoding for photon loss $\gamma$ and measurement error $q$ is compared against the binomial encoding for different values of $N$ and $K$ with each data point being plotted at the effective qubit measurement error rate for that value of $N$ and $K$. Method 2, which involves both local ML QSI and adaptive phase measurement, is used. Thresholds of the binomial codes are now competitive with the trivial encoding at comparable measure error levels.
The final improvement to threshold results we can make is by using Dijkstra decoding to improve MWPM. Dijkstra decoding utilises the continuous variable nature of bosonic codes to weight error paths and improve the accuracy of the matching step of the MWPM algorithm. Fig. 14 shows that using Dijkstra decoding significantly increases threshold values relative to MWPM decoding. Interestingly, we see $N=3$ and $N=2$ outperforming the $N=4$ code. For codes of higher $N$, it is more difficult to distinguish between |±$|\pm\rangle_{L}$ codewords as their angular separation in phase space is less and in addition they have higher photon numbers $\bar{n}$ leading to more photon loss events. Despite the increased loss tolerance of the high $N$ codes it appears that $N=3$ is the optimal $N$ value when subject to these competing effects.
For both plain MWPM and Dijkstra decoding, there appears to be a value of measurement error at ${}1\%$ where thresholds peak. This corresponds to an optimal choice of $K$ for each value of $N$. Higher values of $K$ offer enhanced phase resolution at the price of increased $\bar{n}$ and thus this optimum corresponds to an optimum where better phase resolution due to higher mean photon number is balanced with increased loss due to higher mean photon number. This behaviour is quite distinct from what we saw in figure 10 where there was only a very weak dependence of the threshold on $K$. This is likely because the phase resolution in Method I is dominated by poor performance of heterodyne phase measurement which has phase uncertainty scaling like $1/\bar{n}$ relative to the AHD phase measurement which has uncertainty scaling like $1/\bar{n}^{3/2}$. This opens a window over which it is possible to improve performance by increasing phase resolution at the price of increased $\bar{n}$.

### v2

@@ Abstract
Rotation symmetric bosonic codes are an attractive encoding for qubits into oscillator degrees of freedom, particularly in superconducting qubit experiments. While these codes can tolerate considerable loss and dephasing, they will need to be combined with higher level codes to achieve large-scale devices. We investigate concatenating these codes with the planar code in a measurement-based scheme for fault-tolerant quantum computation. We focus on binomial codes as the base level encoding, and estimate break-even points for such encodings under loss for various types of measurement protocol. These codes are more resistant to photon loss errors, but require both higher mean photon numbers and higher phase resolution for gate operations and measurements. We find that it is necessary to implement adaptive phase measurements, maximum likelihood quantum state inference, and weighted minimum weight decoding to obtain good performance for a planar code using binomial code qubits.
@@ I Introduction
We obtain threshold values for varying orders of discrete rotational symmetry. We find that whilst increasing the order of discrete rotational symmetry generally improves the threshold value, thresholds for binomial encoded qubits tend to be lower than for the trivial Fock space encoding, $|0\rangle_{L}=|0\rangle,|1\rangle_{L}=|1\rangle$, when using the most naive measurement and qubit state inference techniques. By using more sophisticated measurements and state inference methods as well as incorporating the soft information from measurements into decoding we are able to find schemes that beat the trivial encoding when qubit state measurement error for the trivial encoding is not too small.
The point at which the lifetime of a bosonic encoded qubit outperforms the trivial encoding is referred to as break even [30]. The competition between the bosonic codes and the trivial encoding arises because the rotationally symmetric codes achieve tolerance to loss errors at the expense of increased Fock number. In comparison, the trivial encoding has no intrinsic robustness to loss but has a very low average number. We find that optimal measurement and state inference is required for the rotational symmetric codes to outperform the trivial encoding. Finally, we examine the sub-threshold performance of binomial codes and find that below $50\%$ of threshold the loss tolerance of a binomial code is more beneficial and codes with increased rotational symmetry perform better than the trivial code and better than codes with reduced symmetry.
This paper is structured as follows. In Sec. II we review rotation symmetric bosonic codes and measurement based quantum computation. In Sec. III we introduce the measurement models and qubit state inference techniques used to realise the computational scheme. Sec. III.4 then quantifies the performance of these methods. In Sec. III.5 we introduce methods used for the 2D planar code, including an X basis Pauli twirling approximation and modified MWPM decoding algorithm. Sec. III.6 compares the code thresholds obtained using these techniques, as well as examining the subthreshold scaling of the code. Finally, in Sec. IV we review our findings and highlight the main conclusions.
@@ II.1 Rotation Symmetric Bosonic Codes
Indeed, $\hat{R}_{2N}=\hat{Z}_{N}.$ The choice of the logical Z operator determines the form of the logical codewords as finite rotated superpositions of a “primitive” state [25] $|\theta\rangle$, where different types of rotation symmetric codes, such as cat and binomial codes, differ only in the choice of primitive. Explicitly, for $\mu\in\{0,1\}$
@@ II.1.1 RSB $C_{Z}$ Gate
where modes $i,j$ involve RSB qubits with discrete rotational symmetry of order $N,M$ respectively. Loss is the dominant imperfection in our model and if the cross-Kerr gate is slow the effect of loss is exacerbated. Both the strength of the loss and the time required for the gate are captured by the parameter $\gamma=\kappa t_{gate}$ which we will use to characterize the noise level. The gate time is chosen such that $\Omega t_{\rm gate}=\pi/NM$ thus $\gamma$ is proportional to the photon loss rate and inversely proportional to the cross-Kerr nonlinearity. In all that follows we only consider CROT gates between identical encodings, so $N=M.$
Note that low cross-Kerr nonlinearity in practical devices might lead to a different coupling non-linearity being preferred, such as a SNAIL [32]. In this case as well the effect of loss will depend on the per photon loss probability during a gate. So while the details of code performance would differ, we expect qualitatively similar results for such a gate.
@@ II.3 Measurement Based Quantum Computation
In this section we will first describe some preliminary notation used for cluster states, before detailing the construction of 1D cluster states. We will then explain how 1D cluster states can be entangled to realise the 2D surface code as a foliation of the repetition code.
@@ II.3.1 Cluster States
In order to construct the 2D surface code in Fig. 2, we start with parallel 1D cluster states, prepared as before and entangle every odd (primal) qubit of each 1D cluster with the corresponding primal qubit of its neighbouring 1D cluster state, via a dual qubit prepared in the |+ket|$\rangle$| + ⟩ state. Every second qubit of the 1D cluster state is also a dual qubit. The resulting cluster state is depicted in Fig. 4. Measuring every qubit in the X basis realises a measurement of parity checks of primal stabilisers given by a product of $X$ on qubits around a primal plaquette. A primal plaquette stabiliser checks for $Z$ type errors on that plaquette. More explicitly, consider the stabilisers of the cluster state. For a plaquette $p$ of the cluster state, the associated stabiliser is $S_{p}=\prod_{e\in p}X_{e}$, where $e$ denote the primal qubits represented by the edges of the given plaquette as in Fig. 3. To teleport logical information across the 2D foliated state, we measure both primal and dual qubits in the X basis. This also provides stabiliser outcomes, from which we extract the error syndrome.
@@
The model we simulate can either be viewed as the foliated 1D repetition code, or equivalently a single 2D timeslice of plaquette stabilizers of the surface code. The noise model we adopt for this scheme includes photon loss occurring during gates and measurement errors on single qubit X measurements, the latter of which arise naturally from the realistic measurement model employed. In the commonly used phenomenological error models for the surface code, both these noise sources map to gate noise rather than phenomenological measurement noise, so that we are still able to observe an error correction threshold for this scheme. Note however that there is only a single logical operator that can be corrected. Therefore while it is possible to learn a lot about the interplay between loss and measurement errors in this code concatenation scheme, a larger scale simulation corresponding to a $2+1D$ surface code simulation would still be desirable.
@@ III.1.4 Measurement Error
In Figs. 5,6 the data for the measurement error rate as a function of mean photon number is generated numerically. We assume $\gamma=0$ in order to decouple the measurement error from the noise due to photon loss. We then simulate many shots of qubit measurement and calculate the fraction of shots where the logical state as given by the Qubit State Inference, see Sec. III.2, differs from the actual answer found according to Eq. 39. By varying $K$, for fixed $N,$ we vary the mean photon number.
@@
Figure 5: Measurement Error vs $\bar{n}=1/2\cdot KN$ for binomial code with different $N$ values using heterodyne measurement.
@@
Figure 6: Measurement Error vs $\bar{n}=1/2\cdot KN$ for binomial code with different $N$ values using AHD measurement.
It is important to note that the ‘measurement error’ we are referring to is measurement error at the level of the bosonic mode. This manifests as Pauli errors of the binomial code qubits, rather than as errors in the binomial code single qubit $X$ measurement outcomes. Correcting binomial code qubit measurement outcomes would require multiple measurement rounds. Measurement error in our context contributes to Pauli noise independently from noise due to photon loss.
@@ III.4.1 Numerical Performance of Phase Measurements
We performed simulations of the minimal instance of the binomial encoded 1D cluster state scheme with a 3-qubit cluster state. Fig. 8 below quantifies the performance of the different measurement schemes. We see that both canonical phase and AHD measurement schemes significantly outperform heterodyne. Furthermore, there does not appear to be a notable difference between the performance of AHD and canonical phase measurements. This indicates that AHD is a good alternative to the optimal canonical phase measurement. Note that the nonzero infidelity at zero $\gamma$ is due to the measurement error inherent described in Sec. III.1.4. Even in the absence of photon loss, due to the finite mean photon number of the binomial encoded qubits, measurement error will lead to qubit-level errors.
@@ III.5.1 X-Basis Pauli Twirl
III.5.2 Modified MWPM Decoder
@@
During the MWPM algorithm, a complete graph is created from nodes of the syndrome graph. Distances are then calculated between all possible combinations of node pairs. If, for a given node pair, edge weights of the syndrome graph are calculated using soft measurement outcomes, distances cannot be precomputed and must be evaluated using modified MWPM algorithm. This means that for each node pair in the complete graph, modified MWPM algorithm must be run again. In contrast, if edge weights for a node pair are given fixed unit weight, modified MWPM algorithm need only be run a single time. We use a hybrid model in which for nodes separated by Manhattan distances greater than $\log_{2}(L)$, edge weights are calculated using observed hard measurement outcomes, whilst for nodes pairs for which $d_{Manhattan}\leq\log_{2}(L)$, edge weights are calculated using soft measurement outcomes. This provides the advantage of exploiting soft information for a more accurate syndrome decoder, whilst avoiding the long runtime of exhaustively calculating shortest paths between every possible node pair combination.
@@
Finally, we note that we benchmark the performance of the binomial code against the trivial Fock space encoding, as well as binomial codes of different orders of discrete rotational symmetry.
@@ III.5.4 Overview of Numerical Model
Error correction: We begin by measuring logical Z before error correction. We choose one of the smooth boundaries. If an odd number of boundary qubits have errors, the logical Z measurement is $-1$, otherwise it is $+1$. Error correction proceeds as in Sec. II.3. Stabiliser measurements are modelled by, for each plaquette, multiplying together the stored values of all primal qubits comprising the plaquette. If a plaquette has an odd number of qubits with errors, there will be an odd number of $-1$ values and so the result of the multiplication will be $-1$. Otherwise, it will be $+1$. Stabilisers with $-1$ measurement outcomes form the error syndrome. The BlossomV implementation of the MWPM algorithm is then used to pair up nodes in the error syndrome. As we are modelling the planar code, there can be an odd number of $-1$ stabiliser outcomes, so it is also possible to pair syndrome nodes with the boundary. If the modified MWPM algorithm decoder is being used, it is used as described in Sec. III.5.2 to calculate weights between pairs of syndrome nodes to feed into MWPM. Otherwise, each edge of the lattice is assigned weight one. The now paired syndrome nodes are used to determine if a logical error has occurred. We pick the same smooth boundary we used for our initial logical Z measurement. For each of the paired syndrome nodes, we check if the path between them crosses our chosen smooth boundary. If an odd number of crosses occur out of all pairs, the logical Z measurement is $-1$, otherwise it is $+1.$ We compare the logical Z measurement before correction to the new logical Z measurement. If they disagree, a logical error has occurred.
@@ III.6.1 Threshold curves
In this section we will refer to codes using heterodyne phase measurement and the binning algorithm QSI as Method 1 Codes, and codes using AHD measurement and local ML QSI as Method 2 Codes. Figs. 10,12,13,14 show the threshold for different codes as a function of measurement error and $\gamma,$ the photon loss rate. Measurement error is related to mean photon number as per Figs. 5,6, and is determined by the code parameters $N,K$ through $\bar{n}=(1/2)KN.$ The thresholds in the plots are determined for each code (fixed by $K,N$) by sweeping $\gamma$ for different code distances.
Fig. 10 shows the threshold for a range of binomial codes against photon loss $\gamma=\kappa t_{\rm gate}$. Binomial codes with varying values of $N$ are compared to the trivial encoding. The data point for a given value of $N$ and $K$ is plotted at the effective measurement error that is determine for this code in Sec. III. This allows a comparison of the threshold for the trivial encoding with a given qubit measurement error. In this initial example, Method 1 codes are used. The most striking feature of this plot is that the trivial encoding significantly outperforms the binomial codes, for all values of $N.$ Furthermore, increasing $N$ does not have an advantageous effect on the threshold, as might be expected. Finally, varying the binomial code parameter $K,$ and therefore varying the effective measurement error (for a code of fixed $N$) also does not appear to significantly change the threshold.
@@
These effects are strongly suggested by Fig. 11, which plots the same data as Fig. 10, except against $\gamma\bar{n}$ rather than $\gamma$. This effectively normalises the data against mean photon number. We see that binomial codes of higher $N$ do better by this measure, and furthermore that binomial codes outperform the trivial encoding. Thus if we look at the thresholds against the probability of a single photon emission occuring $\gamma\bar{n}$ rather than against the probability for each photon to be lost $\gamma$ then indeed there is increased tolerance to photon loss errors for the higher $N$ codes, although it is the original thresholds measured against $\gamma$ that matter in practice.
@@
Figure 14: Threshold for photon loss errors $\gamma$ with modified MWPM decoding. The trivial encoding for photon loss $\gamma$ and measurement error $q$ is compared against the binomial encoding for different values of $N$ and $K$ with each data point being plotted at the effective qubit measurement error rate for that value of $N$ and $K$. Method 2, which involves both local ML QSI and adaptive phase measurement, is used. Thresholds of the binomial codes are now competitive with the trivial encoding at comparable measure error levels.
The final improvement to threshold results we can make is by using modified MWPM decoding to improve MWPM. Modified MWPM decoding utilises the continuous variable nature of bosonic codes to weight error paths and improve the accuracy of the matching step of the MWPM algorithm. Fig. 14 shows that using modified MWPM decoding significantly increases threshold values relative to MWPM decoding. Interestingly, we see $N=3$ and $N=2$ outperforming the $N=4$ code. For codes of higher $N$, it is more difficult to distinguish between |±$|\pm\rangle_{L}$ codewords as their angular separation in phase space is less and in addition they have higher photon numbers $\bar{n}$ leading to more photon loss events. Despite the increased loss tolerance of the high $N$ codes it appears that $N=3$ is the optimal $N$ value when subject to these competing effects.
As we noted above the measurement errors resulting from imperfect phase measurement depend on the parameters of the binomial code. Higher values of $K$ offer enhanced phase resolution at the price of increased $\bar{n}$ as shown in Figs. 5,6. For this reason the best choice of code balances better phase resolution with increased loss due to higher mean photon number. For both regular and modified MWPM decoding, this optimum appears to be at a value of measurement error ${}1\%$. This corresponds to an optimal choice of $K$ for each value of $N$. This behaviour is quite distinct from what we saw in Fig.  10 where there was only a very weak dependence of the threshold on $K$. This is likely because the phase resolution in Method I is dominated by poor performance of heterodyne phase measurement which has phase uncertainty scaling like $1/\bar{n}$ relative to the AHD phase measurement which has uncertainty scaling like $1/\bar{n}^{3/2}$. This opens a window over which it is possible to improve performance by increasing phase resolution at the price of increased $\bar{n}$.